OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (–1, 3) and (1, 7)?

OpenStudy (anonymous):

Would the answer be, y – 3 = 2(x – 1)

OpenStudy (anonymous):

I was about to write it! yes you're right

OpenStudy (anonymous):

y-3 = 2x -2 y = 2x +1